FEL Upgrade Project Weekly Brief
October 8-12, 2001

We began the first official run week of the last operation’s period for the IR Demo this week.

The FEL was operated for carbon nanotube production, measurements of the optical pulse length, Compton x-ray production and tune-up for the next week’s operations.

We had an extended teleconference with the ONR Contract monitor this week to review the close-out of the Phase 1 contract on the FEL Upgrade project which will be reviewed in detail during the project semiannual review on Nov. 8-9th.  We also discussed a preliminary agenda for the semiannual review.

A budget proposal for FY03-04 was prepared for the Commonwealth of Virginia for support of FEL operations and an addition to the FEL Building to house the Helios synchrotron with additional laboratory and support space.  These proposals are due at the Commonwealth Dept. of Planning and Budget on Monday (Oct. 14th).

We were alerted that the High Energy Laser panel of the Defense Science Board has released their report this week which calls for an increase in DOD spending for laser R&D from $100M/yr to $150M/yr-including work on free electron lasers.

WBS 3 (Beam Physics):
IR Upgrade Driver Design Revision 1.1.2 was completed.  Work started on UV Upgrade Driver Design Revision 1.1.2.  An entertaining feature of energy recovery during multipass operation of the Demo driver was documented as JLAB-TN-01-048, and is available on line at:   http://www.jlab.org/~douglas/FEL/technote/JLABTN01048.pdf

Not only is it possible to simultaneously manage momentum spread on the coasting and energy recovery passes one can also (for a "magic phase") image the injected beam at the dump.

WBS 4 (Injector):
The drawings for the modified cathode shield design were submitted to the shop for manufacturing. Drawings for the shield operating linkage are completed and under review. The gun was put through a heat clean/high voltage cycle to restore the cathode.  It is suspected that during the high voltage processing the conditioning resistor shorted out and a good portion of the conditioning took place without the resistor.  The electrode appears to have some damage which caused us to reduce the operating voltage of the gun to 300kV.  We will be examining the conditioning resistor during a sufficiently long break in machine operations and will be making recommendations for possible reliability improvements for the upgraded injector.

Despite these problems last weekend, the gun has been operating reliably during FEL operations this week.

Gun HVPS - Looking at possibilities of eliminating the transmission line elbow and hangers since this has been a source of arcing during gun conditioning.

WBS 6 (RF):
Zone 3 - Operating the zone into shorts.  Calibration curves have been run on all 8 klystron/ cavity pairs. The final conformal cables need to be installed.

Zone 4 - Two Stepper Cards are on order and due early next month.  Conformal cables need to be installed to make the zone ready for operations.

Quarter HVPS - The design is at the final stage of finding vendors for practical inductors and capacitors.  The rough layout of the HVPS components is in progress.

Quarter Klystrons - A lifting plate for the 100 kW is due next week from our Machine Shop.  The first solenoid is in place now.

WBS 8 (Instrumentation):
The new digital IR cameras are being commissioned. These are used as optical beam position monitors for the FEL transport line.  They have both analog and digital outputs, the power and analog connections have been broken out for immediate use. Digital control cables are on order and will arrive next week. The 30Hz sync and distribution system is operational and ready for testing when the control cable arrive.

The four channel remote relay box is fully operational and documented. It allows the equipment being used in User Lab 3 to be operated from the control room via EPICS.  Response and delay time can then be programmed allowing a more accurate and repeatable measurement.

Bar Codes are being developed for component tracking during the Demo rip-out and subsequent installation of new equipment.  Material and equipment is being inventoried and staged for use and will be included in the data base as the system is developed.

The new Beamviewer control boxes are back from fabrication.  Assembly of the beamviewers and control boxes will begin soon.

The new 32 channel picomotor relay chassis is complete and ready for testing.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
o  Bulk steel, copper and pole tip sheet steel were shipped to Magnet Enterprises International.
o  We  plan to inspect the first coil potting in five weeks.
Injector Dipoles (DU/DV)
o  The drawings continue in final review while we compose the manufacturing specification.
Arc Dipoles (GY, GX, GQ)
o  Design continued on backchecking the detail drawings of the bend and Reverse Bend Dipoles (GX & GQ) with
    the comments and corrections.
o  The Drawings and manufacturing specifications for the 180° Dipole (DY) are ready for signature.
Magnet Measurement
o  Work continued on the dipole and quadrupole measurement stand design.
3 inch quad (QX)
o  New England Technicoil, the core and assembly vendor has completed the first article core.  It will be inspected on
Trim Quad (QT)
o  This magnet remains in a procurement cycle.
Sextupole (SF)
o  Work continued on the magnetic model.
Octupole (OT)
o  DULY Research is working on a preliminary mechanical layout of the magnet.
Corrector Dipoles (DB, DJ)
o  Milhous Co. has not shipped the coils that were due per the latest schedule.  The  20% of the coils manufactured
    report of last week was false.  We are working at getting the vendor to perform to contract.
Beam Line and Vacuum
o  We re-addressed  the features of the Arc Chambers in conjunction with the contract with AES to produce the
    production drawings.
o  The bids for girders and pedestals necessary for the return leg and the hardware for the entire machine were
    received and are in evaluation.
o  Design work continues on the regions between the cryomodules.
o  Work continued on the injector line stands.

WBS 10 (Wiggler):
By rearranging the magnet test setup and using a shorter cable, M. Necaise was able to greatly reduce the noise leading to DC offsets in the magnet measurement setup.  We are now ready to take production measurements.  The downstream vacuum chamber is complete except for the inner flange weld.  The upstream chamber is awaiting final welding after being cleaned.  We expect to have all chambers ready for fit up by the middle of next week.

WBS 11 (Optics):
Procurements continue to be made on optical cavity components. Detailing continues.  Meetings were held with prospective vendors of components for the optical cavity and diagnostics.

The following work was done to support operations:

The protected-silver coated silicon mirror that we installed in the collimator last week only lasted several hours at high power before falling power with a concomitant increase in mode size and lower beam quality indicated that it was heating.  When we pulled the mirror for inspection, we saw that the failure modality was similar to what's been seen earlier (the entire coating is discolored) but the color was different, so we're working with the vendor to determine what's different about their coating.  We assisted the UVa team with installation of their molecular beam apparatus in Lab 6.  We assisted the nanotube team (CWM & NASA) with optics hardware.

The new IR camera installed in one of the OBPM positions is far more sensitive than the older one, so we needed to install additional neutral density filters.  Another near-Brewster windows was installed in User Lab 6.  The FEL autocorrelator was realigned and set up to utilize two-photon absorption, rather than Type I SHG.

We began the week re-establishing beam from the gun after a heat clean cycle.  The high voltage processing went poorly with frequent arcs.  It was later discovered that the conditioning resistor had not been in the circuit, probably due to a short, and we believe this has led to minor damage on the cathode which is presently limiting our operating voltage to 300 kV or below.  We have established operation at this voltage and it is stable but the performance at the shortest wavelengths or on harmonics is an open question for the rest of the run.

We then performed a set of nanotube production experiments producing significant quantities (significant fractions of a gram) of nanotubes for analysis.

We established operation of the FEL autocorrelator and got good measurements of the optical pulse length for comparison with predictions from the spectral width.  We intend to continue these experiments parasitically to get the measurements as a function of FEL detuning.

We continued Thompson X-ray measurements attempting to set up a double scattering measurement of the X-ray signal.

On Friday we performed absorption measurements on optics in our test stand system. These tests were underway as the report was written.